The half-life of the radioactive gas, radon, is 4 days. How much radon will exist after 20 days if the original amount of radon was 200 grams? Use A=200(1/2)^t/4
Answer by stanbon(75887)

The half-life of the radioactive gas, radon, is 4 days.
How much radon will exist after 20 days if the original amount of radon was 200 grams?
Use A(t) = 200(1/2)^(t/4)
----
A(20) = 200(1/2)^5
---
A(20) = 200*(1/32) = 6.25 grams
====================================
Cheers,
Stan H.
==========